# LARES: Latent Reasoning for Sequential Recommendation

### v1

@@ Abstract.
Sequential recommender systems have become increasingly important in real-world applications that model user behavior sequences to predict their preferences. However, existing sequential recommendation methods predominantly rely on non-reasoning paradigms, which may limit the model’s computational capacity and result in suboptimal recommendation performance. To address these limitations, we present LARES, a novel and scalable LAtent REasoning framework for Sequential recommendation that enhances model’s representation capabilities through increasing the computation density of parameters by depth-recurrent latent reasoning. Our proposed approach employs a recurrent architecture that allows flexible expansion of reasoning depth without increasing parameter complexity, thereby effectively capturing dynamic and intricate user interest patterns. A key difference of LARES lies in refining all input tokens at each implicit reasoning step to improve the computation utilization. To fully unlock the model’s reasoning potential, we design a two-phase training strategy: (1) Self-supervised pre-training (SPT) with dual alignment objectives; (2) Reinforcement post-training (RPT). During the first phase, we introduce trajectory-level alignment and step-level alignment objectives, which enable the model to learn recommendation-oriented latent reasoning patterns without requiring supplementary annotated data. The subsequent phase utilizes reinforcement learning (RL) to harness the model’s exploratory ability, further refining its reasoning capabilities. Comprehensive experiments on real-world benchmarks demonstrate our framework’s superior performance. Notably, LARES exhibits seamless compatibility with existing advanced models, further improving their recommendation performance.
@@ 1. Introduction
Recent advances in large reasoning models (DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a) have demonstrated that scaling test-time computation is another effective approach to increasing the utilization of computing power and can significantly enhance the reasoning capabilities of LLMs. This suggests a promising approach to improving model performance by increasing the computation density for each parameter. There are two primary paradigms for test-time scaling in LLMs: one is explicit reasoning (Guo et al., 2024; Shao et al., 2024; Yang et al., 2024, 2025), where models verbalize intermediate reasoning steps (i.e., chain-of-thoughts (CoTs)) by generating meaningful tokens before producing final answers, and the other is latent reasoning (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025), where models perform multi-step implicit reasoning in the latent space without generating explicit tokens. While most LLMs adopt explicit CoT reasoning, this approach faces challenges in recommendation systems. Unlike LLMs operating in dense textual spaces, most sequential recommenders are confined to sparse item ID spaces. This fundamental difference makes it hard to define meaningful reasoning steps like CoTs and provide supervision signals for training (Lightman et al., 2024; Luo et al., 2024). Therefore, we adopt the latent reasoning approach to scaling sequential recommenders. Recent work like ReaRec (Tang et al., 2025) has demonstrated the potential of this paradigm through autoregressive generation of implicit reasoning tokens. Despite its effectiveness, this method remains computationally suboptimal as it only enriches one token per reasoning step.
@@
To tackle this problem, we aim to fully unleash computation by leveraging all input tokens at each reasoning step. Our approach is inspired by the recently proposed depth-recurrent latent reasoning paradigm in LLMs (Geiping et al., 2025), which leverages a middle recurrent architecture to realize test-time scaling by iterating a middle recurrent block.
@@
To this end, we propose LARES, a novel and scalable LAtent REasoning framework for Sequential recommendation that enables flexible test-time scaling by thinking in continuous latent spaces. Our approach adopts a recurrent architecture comprising two key components: a pre-block for initial processing and a core-block for iterative refinement. The core-block supports arbitrary iteration depth, allowing for dynamic computation scaling while improving computational density (i.e., the amount of computation per parameter). To fully unlock the model’s reasoning capabilities, we develop a two-stage training strategy, the self-supervised pre-training (SPT) and reinforcement post-training (RPT). During the SPT stage, we propose the trajectory-level alignment and step-level alignment objectives to equip the model with recommendation-oriented latent reasoning patterns. Specifically, in trajectory-level alignment, we want to achieve mutual enhancement between reasoning actions for the same target. In step-level alignment, we aim to improve the thinking coherence among different intermediate steps. During the RPT stage, we employ reinforcement learning to further refine the model’s reasoning capabilities for recommendation.
@@ 2.1. Overview
Sequential Recommendation. In recommender systems, there are a set of users $\mathcal{U}$ and a set of items ${\mathcal{V}}$. Let $M=|\mathcal{U}|$ and $N=|\mathcal{V}|$ denote the size of the user set and item set. The behavior record for each user $u\in\mathcal{U}$ is defined as an interaction sequence $S_{u}=[v_{1},\dots,v_{n}]$ at the time step $n$, where items are arranged in chronological order. As a typical sequential recommendation setting, traditional sequential recommenders (Zhou et al., 2022; Kang and McAuley, 2018; Zhou et al., 2020) encode the interaction sequence by direct inference to obtain sequential behavior representations. The recommender then predicts the next item the user is most likely to interact with based on the similarities between the encoded user representations and candidate item representations. The objective of next item prediction can be formally written as:
(1) $\max_{\Theta}P(v_{n+1}|\{v_{1},\dots,v_{n}\},\Theta),$
@@
Latent Reasoning For Sequential Recommendation. Traditional sequential recommenders typically adopt a straightforward inference pattern to directly provide recommendations, which struggle with complex recommendation tasks. In contrast, we propose a latent reasoning sequential recommender which is capable of thinking in a continuous space before making recommendations. This multi-step thinking paradigm allows the model to iteratively refine its inference in latent space and obtain more accurate user interests, simulating human-like mental thinking processes when addressing challenging problems. Specifically, the model progressively derives a series of intermediate thoughts, denoted as $T_{u}=[t_{1},...,t_{k}]$ where $k$ is the number of total reasoning steps, and makes final recommendations conditioned on these latent thoughts as well as the behavior sequence. Formally, the objective of latent reasoning sequential recommenders can be formulated as:
(2) $\max_{\Theta}P(v_{n+1},T_{u}|S_{u},\Theta)=P(v_{n+1}|T_{u},S_{u},\Theta)\cdot P(T_{u}|S_{u},\Theta).$
@@
Scaling Latent Reasoning With Recurrent Depth. We present an overview of our proposed framework LARES, as depicted in Figure 1. The architecture adopts a depth-recurrent design consisting of a pre-block for mapping initial features into latent space and a core-block iterable to arbitrary depths, enabling flexible test-time scaling without additional parameters. To facilitate effective latent reasoning, we adopt a two-stage training paradigm for LARES that progressively learns how to think and model genuine user interests in the latent space. This approach draws inspiration from established practices in natural language processing, where pre-training enables LLMs to acquire fundamental knowledge, while reinforcement learning in the post-training phase enhances their reasoning capacity for complex tasks like mathematics, coding and scientific problems, even without supervised fine-tuning (SFT) data (Jaech et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a). While some studies (Dang et al., 2025; Yue et al., 2025) suggest that reinforcement learning may not expand LLMs’ capability boundaries but instead accelerate solution discovery, these findings collectively emphasize the foundational importance of base model ability. Accordingly, our framework first employs self-supervised pre-training to equip LARES with core user interest inference capabilities through iterative latent reasoning, then applies reinforcement post-training to further enhance its reasoning performance by exploring diverse latent thought patterns.
@@
The remainder of this paper is organized as follows: In Section 2.2, we introduce the depth-recurrent architecture of the latent reasoning recommender. Section 2.3 presents the details of self-supervised pre-training, including trajectory-level alignment and step-level alignment, enabling the model to possess the basic ability to infer user interests through a multi-step latent thinking process. In Section 2.4, we describe the process of reinforcement post-training, which leverages reinforcement learning algorithms to encourage the model to explore more thinking paths and further enhance the recommendation performance.
@@ 2.2. Latent Reasoning with A Recurrent Architecture
Our proposed framework, LARES, is a depth-recurrent sequential model consisting of multiple transformer layers, which can be adapted to other advanced sequential recommendation architectures,e.g., FMLPRec (Zhou et al., 2022), BSARec (Shin et al., 2024), and TedRec (Xu et al., 2024), as proved in the experiments in Section 3.4.1. These layers are primarily divided into two groups, the pre-block $\mathcal{A}$ and the core-block $\mathcal{C}$. The pre-block transforms the item embeddings into the latent space and then the core-block performs a multi-step reasoning on the latent representations in a recurrent pattern. Finally, the last latent thought representation is employed for subsequent similarity computation. The most significant difference of LARES lies in the recurrent design of the core block, allowing the model to dive into a deeper thinking of user behaviors without introducing extra parameter burdens. This design is inspired by the findings that the reasoning performance is up to a large effective depth but not necessarily many parameters in natural language processing (Saunshi et al., 2025; Geiping et al., 2025). The advantages of the recurrent architecture are twofold: on the one hand, it enhances the model’s computational expressiveness with no extra parameter memory cost; on the other hand, it enables more flexible inference scaling by controlling the recurrent depth.
Suppose the historical behavior sequence of user $u$ is $S_{u}=[v_{1},\dots,v_{n}]$. Given the item embedding table $\bm{E}\in\mathbb{R}^{N\times d}$, the sequence $S_{u}$ is first transformed into item embeddings $\bm{E}_{u}=[\bm{e}_{1};\dots;\bm{e}_{n}]$ by table lookup, where $d$ is the embedding dimension and [;][;][ ; ] denotes the concatenation operation. First, $\bm{E}_{u}$ along with the position embeddings $\bm{P}\in\mathbb{R}^{n\times d}$ is fed into the pre-block $\mathcal{A}$ to produce the initial latent item representation $\bm{H}\in\mathbb{R}^{t\times d}$, which is written as:
(3) $\bm{H}=\mathcal{A}\left(\bm{E}_{u}+\bm{P}\right).$
Second, the recurrent core-block $C$ performs iterative latent reasoning taking $\bm{H}$ and the latent thought of the previous step $\bm{T}_{i}$ as inputs, formulated as:
@@
(5) $\bm{T}_{0}$ $=\omega*\bm{z},\ \bm{z}\sim\mathcal{N}(0,\sigma_{1}^{2}I),$
where $i$ and $k$ denote the $i$th step and the number of total reasoning steps, $f$ is the aggregation function e.g., addition and concatenation, $\operatorname{LN}[\cdot]$ denotes the layer normalization operation and $\sigma_{1}$ is the standard deviation of the normal distribution for initializing the random state $\bm{T}_{0}$ and $\omega$ is a rescale coefficient. The input $\bm{H}$ to core-block $\mathcal{C}$ in every reasoning step functions as a residual connection to ensure stable gradient backpropagation. Finally, the final latent thought state corresponding to the last item $v_{n+1}$ is regarded as the final user preference representation $\bm{p}\in\mathbb{R}^{d}$ denoted as $\bm{p}=\bm{T}_{k}[-1]$. The output $\bm{p}$ is then used to compute the recommended probabilities of candidate items $\hat{y}_{i}$, and we adopt the widely used cross-entropy loss as the recommendation objective, which is formulated as:
@@ 2.3. Self-Supervised Pre-training for Thinking Adaptation
To ensure flexible scaling of inference-time reasoning, we sample the iteration count $k$ per training step from a log-normal Poisson distribution. For a target mean iteration count $\bar{k}+1$ (where $\bar{k}\in\mathbb{N}$) and variance $\sigma_{2}$, the sampling procedure is defined as follows:
(8) $k$ $\sim\mathcal{P}(e^{\xi})+1,$
where $\mathcal{N}$ and $\mathcal{P}$ denote the normal and Poisson distributions, respectively. In our experiments, we set $\sigma_{2}=0.5$. This distribution predominantly samples values below $\bar{k}$, while retaining a heavy tail that occasionally yields significantly higher iteration counts. To mitigate excessive memory consumption, we truncate $k$ such that $k=\min(k,3\bar{k})$.
@@ 2.3.1. Trajectory-Level Alignment
To further strengthen LARES’s reasoning capability, we introduce trajectory-level alignment to leverage complementary strengths from different reasoning trajectories. During inference, stochastic elements like random initialization and dropout naturally lead to varied reasoning paths across forward passes. We capitalize on this behavior by encouraging knowledge transfer between high-quality reasoning processes. Specifically, we align the final outputs of independent reasoning trajectories for identical input sequences. Following the approach in DuoRec (Qiu et al., 2022), we expand positive instance diversity by treating training sequences sharing the same target item as auxiliary positive views. However, our experiments reveal that aligning two reasoning outcomes with different step lengths adversely impacts model performance. We posit that this occurs because the inconsistency in reasoning steps causes a misalignment between long-chain and short-chain reasoning. Specifically, forcing short-chain reasoning to capture the richer and more complex patterns in longer reasoning processes is inherently challenging. As a result, the model tends to degenerate toward short-chain reasoning, ultimately compromising its effectiveness. To address this, we ensure consistent reasoning steps within each positive pair. Formally, given two positive sequences $S_{u}$ and $\hat{S}_{u}$ and a shared reasoning step $k$, LARES produces final preference representations $\bm{p}=\bm{T}_{k}[-1]$ and $\hat{\bm{p}}=\hat{\bm{T}}_{k}[-1]$. We achieve the alignment between them based on the InfoNCE loss. The trajectory-level alignment objective is formulated as:
@@
where $F(\cdot,\cdot,\cdot)$ denotes InfoNCE loss function, $\mathcal{R}_{\bm{p}}$ and $\mathcal{R}_{\hat{\bm{p}}}$ represent sample sets containing both positive and negative instances.

### v2

@@ Abstract.
Sequential recommender systems have become increasingly important in real-world applications that model user behavior sequences to predict their preferences. However, existing sequential recommendation methods predominantly rely on non-reasoning paradigms, which may limit the model’s computational capacity and result in suboptimal recommendation performance. To address these limitations, we present LARES, a novel and scalable LAtent REasoning framework for Sequential recommendation that enhances model’s representation capabilities through increasing the computation density of parameters by depth-recurrent latent reasoning. Our proposed approach employs a recurrent architecture that allows flexible expansion of reasoning depth without increasing parameter complexity, thereby effectively capturing dynamic and intricate user interest patterns. A key difference of LARES lies in refining all input tokens at each implicit reasoning step to improve the computation utilization. To fully unlock the model’s reasoning potential, we design a two-phase training strategy: (1) Self-supervised pre-training (SPT) with dual alignment objectives; (2) Reinforcement post-training (RPT). During the first phase, we introduce trajectory-level alignment and step-level alignment objectives, which enable the model to learn recommendation-oriented latent reasoning patterns without requiring supplementary annotated data. The subsequent phase utilizes reinforcement learning (RL) to harness the model’s exploratory ability, further refining its reasoning capabilities. Comprehensive experiments on real-world benchmarks demonstrate our framework’s superior performance. Notably, LARES exhibits seamless compatibility with existing advanced models, further improving their recommendation performance. Our code is available at https://anonymous.4open.science/r/LARES-E458/.
@@
GSAI is the abbreviation of Gaoling School for Artificial Intelligence.
@@ 1. Introduction
Recent advances in large reasoning models (DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a) have demonstrated that scaling test-time computation is another effective approach to increasing the utilization of computing power and can significantly enhance the reasoning capabilities of LLMs. This suggests a promising approach to improving model performance by increasing the computation density for each parameter. There are two primary paradigms for test-time scaling in LLMs: one is explicit reasoning (Guo et al., 2024; Shao et al., 2024; Yang et al., 2024, 2025), where models verbalize intermediate reasoning steps (i.e., chain-of-thoughts (CoTs)) by generating meaningful tokens before producing final answers, and the other is latent reasoning (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025), where models perform multi-step implicit reasoning in the latent space without generating explicit reasoning tokens. While most LLMs adopt explicit CoT reasoning, this approach faces challenges in recommendation systems. Unlike LLMs operating in dense textual spaces, most sequential recommenders are confined to sparse item ID spaces. This fundamental difference makes it hard to define meaningful reasoning steps like CoTs and provide supervision signals for training (Lightman et al., 2024; Luo et al., 2024). Therefore, we adopt the latent reasoning approach to scaling sequential recommenders. Recent work like ReaRec (Tang et al., 2025) has demonstrated the potential of this paradigm through autoregressive generation of implicit reasoning tokens. Despite its effectiveness, this method remains computationally suboptimal as it only enriches one token per reasoning step.
@@
To tackle this problem, we aim to fully unleash computation by leveraging all input tokens at each reasoning step. Our approach is inspired by the recently proposed depth-recurrent model architecture for latent reasoning (Geiping et al., 2025), which designs a recurrent block consisting of Transformer layers. It repeatedly applies the recurrent block to update the hidden states of all the tokens in the latent space for test-time compute scaling. For the scaling of sequential recommenders, since we also aim to unleash the computation of each item token in the latent space, it is feasible to develop a depth-recurrent architecture over existing recommendation models.
@@
To this end, we propose LARES, a novel and scalable LAtent REasoning approach for Sequential recommendation that enables flexible test-time scaling by thinking in continuous latent spaces. Our approach adopts a recurrent architecture comprising two key components: a pre-block for initial processing and a core-block for iterative refinement. The core-block supports arbitrary iteration depth, allowing for dynamic computation scaling while improving computational density (i.e., the amount of computation per parameter). To fully unlock the model’s reasoning capabilities, we develop a two-stage training strategy, the self-supervised pre-training (SPT) and reinforcement post-training (RPT). During the SPT stage, we propose the trajectory-level alignment and step-level alignment objectives to equip the model with recommendation-oriented latent reasoning patterns. Specifically, in trajectory-level alignment, we want to achieve knowledge transfer between high-quality reasoning processes. In step-level alignment, we aim to improve the thinking coherence among different intermediate steps. During the RPT stage, we employ reinforcement learning to further refine the model’s reasoning capabilities for recommendation.
@@ 2.1. Overview
Sequential Recommendation. In recommender systems, there are a set of users $\mathcal{U}$ and a set of items ${\mathcal{V}}$. Let $M=|\mathcal{U}|$ and $N=|\mathcal{V}|$ denote the size of the user set and item set. The behavior record for each user $u\in\mathcal{U}$ is defined as an interaction sequence $S_{u}=[v_{1},\dots,v_{n}]$ at the time step $n$, where items are arranged in chronological order. As a typical sequential recommendation setting, traditional sequential recommenders (Kang and McAuley, 2018; Zhou et al., 2020; Sun et al., 2019) encode the interaction sequence by direct inference to obtain sequential behavior representations. The recommender then predicts the next item the user is most likely to interact with based on the similarities between the encoded user representations and candidate item representations. The objective of next item prediction can be formally written as:
(1) $\max_{\Theta}P(v_{n+1}|S_{u};\Theta),$
@@
Latent Reasoning For Sequential Recommendation. Traditional sequential recommenders typically adopt a straightforward inference pattern to directly provide recommendations, which struggle with complex recommendation tasks. In contrast, recent studies (Tang et al., 2025) propose a latent reasoning sequential recommender which is capable of thinking in a continuous space before making recommendations. This multi-step thinking paradigm allows the model to iteratively refine its inference in latent space and obtain more accurate user interests, simulating human-like mental thinking processes when addressing challenging problems. Generally, the model progressively derives a series of intermediate thoughts, denoted as $T_{u}$ and makes final recommendations conditioned on these latent thoughts as well as the behavior sequence. Formally, the objective of latent reasoning sequential recommenders can be formulated as:
(2) $\max_{\Theta}P(v_{n+1}|S_{u};\Theta)=P(v_{n+1}|T_{u},S_{u};\Theta)\cdot P(T_{u}|S_{u};\Theta).$
Existing work (Tang et al., 2025) models latent reasoning $P(T_{u}|S_{u};\Theta)$ as an autoregressive generation process, where a new state is generated and appended to the input at each reasoning step, which can be represented as follows:
(3) $P(T_{u}|S_{u};\Theta)=\prod_{i=1}^{k}P(t_{i}|S_{u},t_{<i};\Theta),$
where $k$ is the number of reasoning steps and $t_{i}$ and $t_{<i}$ denote the generated thoughts at the $i$-th step and preceding the $i$-th step, respectively.
Scaling Latent Reasoning With Recurrent Depth. To fully unleash the computing power of deep thinking in sequential recommendation, we propose LARES, a novel latent reasoning scaling paradigm with recurrent depth where all input tokens are refined at each reasoning step instead of only generating one new token. The reasoning process of LARES is denoted as follows:
(4) $P(T_{u}|S_{u};\Theta)=\prod_{i=1}^{k}P(S^{i}_{u}|S^{i-1}_{u},S^{0}_{u};\Theta),$
@@
where $k$ is the number of reasoning steps (i.e., recurrent depth), $S^{i}_{u}$ is the thought at the $i$-th step, and $S^{0}_{u}$ corresponds to the initial input. The overall framework of LARES is depicted in Figure 1. It adopts a depth-recurrent design consisting of a pre-block for mapping initial features into latent space and a core-block iterable to arbitrary depths, enabling flexible test-time scaling without additional parameters. To facilitate effective latent reasoning with only the outcome label (i.e., the next item), we propose a two-stage training approach for LARES: self-supervised pre-training for thinking adaptation and reinforcement post-training for thinking exploration. This approach draws inspiration from established practices in LLMs, where pre-training enables LLMs to acquire fundamental knowledge, while reinforcement learning incentivizes the reasoning capability for complex reasoning tasks (Jaech et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a). Accordingly, our framework first employs self-supervised pre-training to equip LARES with the core capability of user interest modeling through iterative latent reasoning, then applies reinforcement post-training to further incentivize its reasoning capability by exploring diverse latent thought patterns.
@@
The remainder of this paper is organized as follows: In Section 2.2, we introduce the depth-recurrent architecture of the latent reasoning recommender. Section 2.3 presents the details of self-supervised pretraining, including trajectory-level alignment and step-level alignment, enabling the model to possess the basic ability to infer user interests through a multi-step latent thinking process. In Section 2.4, we describe the process of reinforcement post-training, which leverages reinforcement learning algorithms to encourage the model to explore diverse thinking paths and further enhance the recommendation performance.
@@ 2.2. Latent Reasoning with A Recurrent Architecture
Our proposed framework, LARES, is a depth-recurrent sequential model consisting of multiple Transformer layers, which can be adapted to other advanced sequential recommendation architectures, e.g., FMLPRec (Zhou et al., 2022), BSARec (Shin et al., 2024), and TedRec (Xu et al., 2024), as proved in the experiments in Section 3.4.1. In LARES, there are two blocks, i.e., the pre-block $\mathcal{A}$ and the core-block $\mathcal{C}$. The pre-block first transforms the item embeddings into the latent space, and then the core-block performs a multi-step reasoning on the latent representations in a recurrent pattern. Finally, the last latent thought representation is employed for subsequent item prediction. The most significant design of LARES lies in the recurrence of the core block, which allows the model to dive into a deeper thinking of user behaviors without introducing extra parameter burdens. This design is inspired by the findings that the reasoning performance is up to a large effective depth but not necessarily many parameters (Saunshi et al., 2025; Geiping et al., 2025). The advantages of the recurrent architecture are twofold: on the one hand, it enhances the model’s computational expressiveness with no extra parameter memory cost; on the other hand, it enables more flexible inference scaling by controlling the recurrent depth.
Suppose the historical behavior sequence of user $u$ is $S_{u}=[v_{1},\dots,v_{n}]$. Given the item embedding table $\bm{E}\in\mathbb{R}^{N\times d}$, the sequence $S_{u}$ is first transformed into item embeddings $\bm{E}_{u}=[\bm{e}_{1};\dots;\bm{e}_{n}]$ by table lookup, where $d$ is the embedding dimension and [;][;][ ; ] denotes the concatenation operation. First, $\bm{E}_{u}$ along with the position embeddings $\bm{E}_{p}\in\mathbb{R}^{n\times d}$ is fed into the pre-block $\mathcal{A}$ to produce the initial latent item representation $\bm{H}\in\mathbb{R}^{t\times d}$, which is written as:
(5) $\bm{H}=\mathcal{A}\left(\bm{E}_{u}+\bm{E}_{p}\right).$
Then, the recurrent core-block $\mathcal{C}$ performs iterative latent reasoning, taking $\bm{H}$ and the latent thought of the previous step $\bm{T}_{i-1}$ as inputs, formulated as:
@@
(7) $\bm{T}_{0}$ $\sim\mathcal{N}(0,\sigma_{1}^{2}I),$
where $i$ and $k$ denote the $i$th step and the number of total reasoning steps, $f$ is the aggregation function e.g., addition and concatenation, $\operatorname{LN}[\cdot]$ denotes the layer normalization operation and $\sigma_{1}$ is the standard deviation of the normal distribution for initializing the random state $\bm{T}_{0}$. The input $\bm{H}$ to core-block $\mathcal{C}$ in every reasoning step functions as a residual connection to ensure stable gradient backpropagation. To allow flexible scaling of inference-time reasoning, during training, we sample the iteration count $k$ per training step from a log-normal Poisson distribution. For a target mean iteration count $\bar{k}+1$ (where $\bar{k}\in\mathbb{N}$) and variance $\sigma_{2}$, the sampling procedure is defined as follows:
(9) $k$ $\sim\mathcal{P}(e^{\xi})+1,$
where $\mathcal{N}$ and $\mathcal{P}$ denote the normal and Poisson distributions, respectively. In our experiments, we set $\sigma_{2}=0.5$. This distribution predominantly samples values below $\bar{k}$, while retaining a heavy tail that occasionally yields significantly higher iteration counts. To mitigate excessive memory consumption, we truncate $k$ such that $k=\min(k,3\bar{k})$. In inference, the recurrent depth is set to $\bar{k}+1$ for all test samples.
Finally, the final latent thought state corresponding to the last item $v_{n+1}$ is regarded as the final user preference representation $\bm{p}\in\mathbb{R}^{d}$ denoted as $\bm{p}=\bm{T}_{k}[-1]$. The output $\bm{p}$ is then used to compute the recommended probabilities of candidate items $\hat{y}_{i}$, and we adopt the widely used cross-entropy loss as the recommendation objective, which is formulated as:
@@ 2.3.1. Trajectory-Level Alignment
To strengthen LARES’s reasoning capability, we introduce trajectory-level alignment to leverage complementary strengths from different reasoning trajectories. We define different trajectories from two aspects: On the one hand, stochastic elements like random initialization and dropout naturally lead to varied reasoning paths across forward passes for the identical input sequences; On the other hand, the trajectory of different sequences sharing the same target item can also be regarded as positive views (Qiu et al., 2022). Specifically, we align the final outputs of independent reasoning trajectories between positive pairs. However, our experiments reveal that aligning two reasoning outcomes with different step lengths adversely impacts model performance. We posit that this occurs because the inconsistency in reasoning steps causes a misalignment between long-chain and short-chain reasoning. Specifically, forcing short-chain reasoning to capture the richer and more complex patterns in longer reasoning processes is inherently challenging. As a result, the model tends to degenerate toward short-chain reasoning, ultimately compromising its effectiveness. To address this, we ensure consistent reasoning steps within each positive pair. Formally, given two positive sequences $S_{u}$ and $\hat{S}_{u}$ and a shared reasoning step $k$, LARES produces final preference representations $\bm{p}=\bm{T}_{k}[-1]$ and $\hat{\bm{p}}=\hat{\bm{T}}_{k}[-1]$. We achieve the alignment between them based on the InfoNCE loss. The trajectory-level alignment objective is formulated as:
@@
where $F(\cdot,\cdot,\cdot)$ denotes the InfoNCE loss function, $s(\cdot,\cdot)$ is a similarity metric (e.g., cosine or dot product), $\mathcal{R}_{\bm{p}}$ and $\mathcal{R}_{\hat{\bm{p}}}$ represent sample sets containing both positive and negative instances, and $\tau$ is a temperature coefficient controlling the distribution sharpness.